OpenStudy (anonymous):

Help me write a real world scenario for the following equation... 2x + y = 10 3x + 4y = 25 Solve this with me and help me explain what the results mean according to the scenario we will create together :D

OpenStudy (anonymous):

I am Russian

OpenStudy (anonymous):

u dnt know how to do this

OpenStudy (anonymous):

There are two methods to solve this system of equations. This is one of them ! From the 1st equation we can say that : \[y=10-2x\] So the 2nd equation becomes like this : \[3x + 4(10-2x) = 25 \] So : \[3x+40-8x=25\] So : \[-5x=25-40\] So : \[-5x= -15\] so : \[x=\frac{-15}{-5}=3\] And we return to the equation of y : \[y=10-2x=10-2\times3=4\]

OpenStudy (anonymous):

Could you please help me to come up with real world situation for this problem please?

OpenStudy (anonymous):

yes